### v1

@@ Abstract
We explore the dynamics of pure scalar fields rolling on an exponential potential in the absence of any additional background fluid and demonstrate the existence of self-tracking solutions in which the self-perturbations of the scalar field act as an effective radiation background. The validity of these solutions is demonstrated through both analytic techniques and numerical simulations using ${\mathcal{C}}$${\mathcal{L}}$attice . We discuss applications to string cosmologies with significant trans-Planckian field excursions between inflation and BBN, including the required initial level of scalar perturbations to avoid overshoot.
@@ 1 Introduction
Building on the seminal work of Ferreira:1997hj ; Copeland:1997et , there exists considerable recent literature Gouttenoire:2021jhk ; Conlon:2022pnx ; Revello:2023hro ; Shiu:2023fhb ; Shiu:2023nph ; Seo:2024qzf ; Apers:2024ffe ; Shiu:2024sbe ; Andriot:2024jsh ; Conlon:2024uob ; Revello:2024gwa ; Cicoli:2024yqh ; Casas:2024oak ; Apers:2022cyl ; Apers:2024dtn ; Eroncel:2025bcb ; Brunelli:2025ems ; Andriot:2025cyi ; Ghoshal:2025tlk ; SanchezGonzalez:2025uco (for older work in a string context see Barreiro:1998aj ; Conlon:2008cj ; Cicoli:2015wja ) which explores cosmologies combining a rolling scalar field and a background fluid that may consist of matter, radiation or even more exotic objects such as cosmic strings. In such cases, the final attractor for these cosmologies is often a tracker solution with fixed proportions of the energy density in the background fluid, the kinetic energy of the scalar field and the potential energy of the scalar field.
@@
The paper is organised as follows. Following the introduction (section 1), we give an analytic argument as to why the self-perturbations of a scalar field on an exponential potential can, effectively, act as a radiation background, and can give rise to a tracker solution (section 2). We then describe the results of numerical simulations using ${\mathcal{C}}$${\mathcal{L}}$attice that confirm the results of the analytic treatment (section 3). We then discuss physics applications including the possibility that such a self-tracker can address the overshooting problem of string compactifications.
@@ 2.1 Perturbations in a Scalar Field
where reality of the scalar field implies that $\delta\phi_{\boldsymbol{k}}^{*}=\delta\phi_{-\boldsymbol{k}}$. Additionally, any spatial averages over sufficiently large scales result in odd powers of the perturbations vanishing $\langle\delta\phi^{2n+1}\rangle=0$. If the potential has an exponential form $V=V_{0}e^{-\alpha\phi/M_{P}}$, it can be expanded out as
$V(\phi(t,\boldsymbol{x}))=V_{0}e^{-\alpha\bar{\phi}(t)/M_{P}-\alpha\delta\phi/M_{P}}=\bar{V}(1-\tfrac{\alpha}{M_{P}}\delta\phi+\tfrac{\alpha^{2}}{2M_{P}^{2}}\delta\phi^{2})+\mathcal{O}(\delta\phi^{3}),$ (2.4)
provided that the field perturbations are small, $\alpha\delta\phi\ll M_{P}$. Here $\bar{V}=V(\bar{\phi})$ is defined as the potential of the background field. The density and pressure of the scalar field can in turn be decomposed as
$\begin{split}\rho_{\phi}&=\bigg{(}\frac{1}{2}\bar{\pi}^{2}+\bar{V}\bigg{)}+\bigg{(}[\bar{\pi}\delta\pi-\tfrac{\alpha}{M_{P}}\bar{V}\delta\phi]+\frac{1}{2}\delta\pi^{2}+\frac{1}{2a^{2}}(\nabla\delta\phi)^{2}+\frac{1}{2}\bar{m}^{2}\delta\phi^{2}+\mathcal{O}(\delta\phi^{3})\bigg{)},\\ P_{\phi}&=\bigg{(}\frac{1}{2}\bar{\pi}^{2}-\bar{V}\bigg{)}+\bigg{(}[\bar{\pi}\delta\pi+\tfrac{\alpha}{M_{P}}\bar{V}\delta\phi]+\frac{1}{2}\delta\pi^{2}-\frac{1}{6a^{2}}(\nabla\delta\phi)^{2}-\frac{1}{2}\bar{m}^{2}\delta\phi^{2}+\mathcal{O}(\delta\phi^{3})\bigg{)},\end{split}$ (2.5)
where $\bar{m}^{2}=\alpha^{2}\bar{V}/M_{P}^{2}$ is the effective mass of the perturbations and $\pi=\dot{\phi}$ (which is also decomposed into its homogeneous and inhomogeneous parts). Spatially averaging these gives
@@
$\bigg{(}\ddot{\bar{\phi}}+3H\dot{\bar{\phi}}-\frac{\alpha}{M_{P}}\bar{V}\bigg{)}+\bigg{(}\delta\ddot{\phi}+3H\delta\dot{\phi}-\frac{1}{a^{2}}\nabla^{2}\delta\phi+\bar{m}^{2}\delta\phi\bigg{)}+\mathcal{O}(\delta\phi^{2})=0.$ (2.8)
@@
$\ddot{\bar{\phi}}+\frac{3c}{t}\dot{\bar{\phi}}-\tfrac{\alpha V_{0}}{M_{P}}e^{-\alpha\bar{\phi}/M_{P}}=0.$ (2.9)
@@
$\bar{\phi}(t)=-\frac{M_{P}}{\alpha}\ln\bigg{(}\frac{6c-2}{\alpha^{2}V_{0}M_{P}^{-2}}\bigg{)}+\frac{2M_{P}}{\alpha}\ln t,$ (2.10)
@@
This solution (2.10) is a stable fixed point. To see this, consider the perturbing around it $\bar{\phi}(t)\rightarrow\bar{\phi}(t)+\delta\bar{\phi}(t)$. To leading order in the expansion of the exponential, the solution to the linearised version of Eq. (2.9) is
@@
$\delta\phi_{\boldsymbol{k}}(t)=c_{1}^{\prime}t^{-(3c-1)/2}J_{\alpha}\bigg{[}\frac{ka^{-1}}{1-c}t\bigg{]}+c_{2}^{\prime}t^{-(3c-1)/2}J_{-\alpha}\bigg{[}\frac{ka^{-1}}{1-c}t\bigg{]},$ (2.17)
where $\alpha=(3c-1)/(2-2c)$ and $c_{1}^{\prime}$ and $c_{2}^{\prime}$ are coefficients fixed by the initial conditions. Using the asymptotic form of the Bessel functions for large real arguments this result can be simplified at either late times or large $k$ to
@@
The perturbations $\delta\phi$ can be treated as a separate radiation fluid distinct from the spatially-averaged scalar field $\bar{\phi}$. This is because at leading order, the only direct coupling between the two in the scalar field equation of motion occurs through the mass term $m^{2}(\bar{\phi})\delta\phi=\alpha^{2}M_{P}^{-2}V(\bar{\phi})\delta\phi$. As long as this term is negligible, as occurs for sub-horizon modes, then the two decouple. Importantly, this is the case irrespective of which component dominates the energy density of the scalar field. In particular, a consistent expansion of the exponential in Eq. (2.4) only requires that the inhomogeneities are small relative to the Planck scale $\alpha\delta\phi\ll M_{P}$, but they need not be small relative to the background scalar field or its time derivative.
@@
For sufficiently steep exponential potentials with $\alpha>\sqrt{2}$, this cosmology has a well-known late-time solution given by the radiation tracker for which $c=1/2$. In this case, such a tracker solution can be made explicit by defining the kinetic and potential energy fraction of the full scalar field as Copeland:1997et
@@
$x=\langle\tilde{x}\rangle=\frac{\bar{\pi}}{\sqrt{6}HM_{P}},\ \ \ \ \ \ y=\sqrt{\frac{V(\bar{\phi})}{3}}\frac{1}{HM_{P}}=\langle\tilde{y}\rangle[1+\tfrac{\alpha^{2}}{8M_{P}^{2}}\langle\delta\phi^{2}\rangle+\mathcal{O}(\langle\delta\phi^{4}\rangle)].$ (2.21)
@@
$x^{\prime}(N)$ $=-3x+\alpha\sqrt{\frac{3}{2}}y^{2}+2x[1+\tfrac{1}{2}x^{2}-y^{2}],$ (2.22)
$y^{\prime}(N)$ $=-\alpha\sqrt{\frac{3}{2}}xy+2y[1+\tfrac{1}{2}x^{2}-y^{2}],$ (2.23)
@@
where primes represent derivatives with respect to $N$. For sufficiently steep exponential potentials with $\alpha>\sqrt{2}$ this has the standard attractor radiation tracker solution
$x^{2}=\frac{8}{3\alpha^{2}},\ \ \ \ \ \ \ \ y^{2}=\frac{4}{3\alpha^{2}}.$ (2.26)
@@ 2.2 Non-linearities and Non-Exponential Potentials
It is useful to contextualize the case of exponential potentials by considering the case of more general potentials $V(\phi)$ (some of these result have recently been pointed out by Eroncel:2025qlk ). In that case the scalar field equation of motion can be written as
@@
[…]
@@
In this work, we have primarily focused on the case of exponential potentials, where the potential factorizes into two exponentials $e^{-\alpha\bar{\phi}/M_{P}}e^{-\alpha\delta\phi/M_{P}}$. The equation for the perturbations can then be expanded out as long as $\alpha\delta\phi\ll M_{P}$, giving a linear equation, with higher order terms suppressed by factors of $\alpha\delta\phi/M_{P}$. Additionally, the approximation becomes more accurate with time as long as the amplitude of the perturbations decreases with time. This is true irrespective of whether $\delta\phi$ dominates the energy density of the universe or not. For example, the transition from a kinating epoch, where the energy density is dominated by the kinetic energy in the background field $\dot{\bar{\phi}}^{2}/2\gg V(\phi)$, to one where the universe is dominated by the perturbations, is perfectly linear in terms of the equation of motion for the perturbations $\delta\phi$.
As well as through the scalar potential, another way in which non-linearities can arise is if there is significant metric backreaction. This can occur when the energy density in some region is substantially larger than the background $\delta\rho/\bar{\rho}\gtrsim 1$ for long enough that the region decouples from the Hubble flow Aurrekoetxea:2024mdy . In that case a full numerical relativistic analysis is required, with these regions possibly giving rise to black holes. The case of a massless scalar field with an initial Guassian profile has been studied in numerical relativity Yoo:2018pda . This system collapses to a black hole for sufficiently large initial amplitudes, but would otherwise transition into a radiation dominated universe if collapse is avoided, which corresponds in our case to perturbation domination with a negligible potential.
@@ 2.3 Perturbation-Dominated Eras
$\frac{1}{2}\bar{m}^{2}\langle\delta\phi^{2}\rangle=\bar{V}\frac{\alpha^{2}}{M_{P}^{2}}\int\frac{d^{3}\boldsymbol{k}}{(2\pi)^{3}}|\delta\phi_{\boldsymbol{k}}|^{2}\leq\bar{V}\bigg{(}\frac{\alpha}{M_{P}}\int\frac{d^{3}\boldsymbol{k}}{(2\pi)^{3}}|\delta\phi_{\boldsymbol{k}}|\bigg{)}^{2}\ll\bar{V},$ (2.29)
@@
$\frac{\alpha}{M_{P}}\int\frac{d^{3}\boldsymbol{k}}{(2\pi)^{3}}|\delta\phi_{\boldsymbol{k}}|\ll 1,$ (2.30)
which is a marginally stronger requirement than $\alpha\delta\phi/M_{P}\ll 1$, needed to be able to expand out the exponential to leading order. Since perturbation domination requires the energy density contribution from the perturbations to be much greater than that of the background potential $\bar{V}$, perturbation domination with small amplitudes cannot behave like a matter dominated era. [^3] For other potentials, this need not be the case. For example, for quadratic potentials where the background scalar field is at its minimum $\bar{V}=0$, one can have a matter dominated era driven by the energy density coming from the perturbations around the minimum.
@@
$3H|\delta\dot{\phi}_{\boldsymbol{k}}|\gg 3H\bar{m}|\delta\phi_{\boldsymbol{k}}|\gg\frac{1}{\alpha}\bar{m}^{2}|\delta\phi_{\boldsymbol{k}}|\sim\bar{m}^{2}|\delta\phi_{\boldsymbol{k}}|,$ (2.33)
where for the second inequality we have used that during perturbation domination $H^{2}\gg\bar{V}/M_{P}^{2}=\bar{m}^{2}/\alpha^{2}$. Therefore, for $\alpha\sim\mathcal{O}(1)$, the mass term is much smaller than the Hubble term. We have thus shown that when the mass contribution to the energy density is much smaller than the gradient/kinetic terms, then the gradient/Hubble terms dominate over the mass term in the equation of motion, respectively. The equation therefore corresponds to that of a massless scalar field, implying that during perturbation domination with subhorizon modes, the universe behaves like a radiation era.
@@
$\delta\ddot{\phi}_{\boldsymbol{k}}+3H\delta\dot{\phi}_{\boldsymbol{k}}+\bigg{(}\frac{k^{2}}{a^{2}}+\frac{\alpha^{2}\bar{V}}{M_{P}^{2}}\bigg{)}\delta\phi_{\boldsymbol{k}}=0,$ (2.34)
which for $k/a<\alpha\sqrt{\bar{V}}/M_{P}$ behaves like a massive scalar field. In a cosmology with $H=c/t$, this can be solved to give
@@
where $\beta=(3c-1)/2$, and in the second step we used the asymptotic form of the Bessel functions for large $mt$. This solution behaves like a matter field with mass $m=\alpha\sqrt{\bar{V}}/M_{P}$. However, so long as the background scalar field can start rolling down the exponential, then such matter behavior would be short-lived.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Building on the seminal work of Ferreira:1997hj ; Copeland:1997et , there exists considerable recent literature Gouttenoire:2021jhk ; Conlon:2022pnx ; Revello:2023hro ; Shiu:2023fhb ; Shiu:2023nph ; Seo:2024qzf ; Apers:2024ffe ; Shiu:2024sbe ; Andriot:2024jsh ; Conlon:2024uob ; Revello:2024gwa ; Cicoli:2024yqh ; Casas:2024oak ; Apers:2022cyl ; Apers:2024dtn ; Eroncel:2025bcb ; Brunelli:2025ems ; Andriot:2025cyi ; Bedroya:2025ris ; Ghoshal:2025tlk ; SanchezGonzalez:2025uco (for older work in a string context see Barreiro:1998aj ; Conlon:2008cj ; Cicoli:2015wja ) which explores cosmologies combining a rolling scalar field and a background fluid that may consist of matter, radiation or even more exotic objects such as cosmic strings. In such cases, the final attractor for these cosmologies is often a tracker solution with fixed proportions of the energy density in the background fluid, the kinetic energy of the scalar field and the potential energy of the scalar field.
@@
[…]
@@ 2.1 Perturbations in a Scalar Field
where reality of the scalar field implies that $\delta\phi_{\boldsymbol{k}}^{*}=\delta\phi_{-\boldsymbol{k}}$. Additionally, any spatial averages over sufficiently large scales result in odd powers of the perturbations vanishing $\langle\delta\phi^{2n+1}\rangle=0$. If the potential has an exponential form $V=V_{0}e^{-\lambda\phi/M_{P}}$, it can be expanded out as
$V(\phi(t,\boldsymbol{x}))=V_{0}e^{-\lambda\bar{\phi}(t)/M_{P}-\lambda\delta\phi/M_{P}}=\bar{V}(1-\tfrac{\lambda}{M_{P}}\delta\phi+\tfrac{\lambda^{2}}{2M_{P}^{2}}\delta\phi^{2})+\mathcal{O}(\delta\phi^{3}),$ (2.4)
provided that the field perturbations are small, $\lambda\delta\phi\ll M_{P}$. Here $\bar{V}=V(\bar{\phi})$ is defined as the potential of the background field. The density and pressure of the scalar field can in turn be decomposed as
$\begin{split}\rho_{\phi}&=\bigg{(}\frac{1}{2}\bar{\pi}^{2}+\bar{V}\bigg{)}+\bigg{(}[\bar{\pi}\delta\pi-\tfrac{\lambda}{M_{P}}\bar{V}\delta\phi]+\frac{1}{2}\delta\pi^{2}+\frac{1}{2a^{2}}(\nabla\delta\phi)^{2}+\frac{1}{2}\bar{m}^{2}\delta\phi^{2}+\mathcal{O}(\delta\phi^{3})\bigg{)},\\ P_{\phi}&=\bigg{(}\frac{1}{2}\bar{\pi}^{2}-\bar{V}\bigg{)}+\bigg{(}[\bar{\pi}\delta\pi+\tfrac{\lambda}{M_{P}}\bar{V}\delta\phi]+\frac{1}{2}\delta\pi^{2}-\frac{1}{6a^{2}}(\nabla\delta\phi)^{2}-\frac{1}{2}\bar{m}^{2}\delta\phi^{2}+\mathcal{O}(\delta\phi^{3})\bigg{)},\end{split}$ (2.5)
where $\bar{m}^{2}=\lambda^{2}\bar{V}/M_{P}^{2}$ is the effective mass of the perturbations and $\pi=\dot{\phi}$ (which is also decomposed into its homogeneous and inhomogeneous parts). Spatially averaging these gives
@@
$\bigg{(}\ddot{\bar{\phi}}+3H\dot{\bar{\phi}}-\frac{\lambda}{M_{P}}\bar{V}\bigg{)}+\bigg{(}\delta\ddot{\phi}+3H\delta\dot{\phi}-\frac{1}{a^{2}}\nabla^{2}\delta\phi+\bar{m}^{2}\delta\phi\bigg{)}+\mathcal{O}(\delta\phi^{2})=0.$ (2.8)
@@
$\ddot{\bar{\phi}}+\frac{3c}{t}\dot{\bar{\phi}}-\tfrac{\lambda V_{0}}{M_{P}}e^{-\lambda\bar{\phi}/M_{P}}=0.$ (2.9)
@@
$\bar{\phi}(t)=-\frac{M_{P}}{\lambda}\ln\bigg{(}\frac{6c-2}{\lambda^{2}V_{0}M_{P}^{-2}}\bigg{)}+\frac{2M_{P}}{\lambda}\ln t,$ (2.10)
@@
This solution (2.10) is a stable fixed point. To see this, consider perturbing around it, $\bar{\phi}(t)\rightarrow\bar{\phi}(t)+\delta\bar{\phi}(t)$. To leading order in the expansion of the exponential, the solution to the linearised version of Eq. (2.9) is
@@
$\delta\phi_{\boldsymbol{k}}(t)=c_{1}^{\prime}t^{-(3c-1)/2}J_{\beta}\bigg{[}\frac{ka^{-1}}{1-c}t\bigg{]}+c_{2}^{\prime}t^{-(3c-1)/2}J_{-\beta}\bigg{[}\frac{ka^{-1}}{1-c}t\bigg{]},$ (2.17)
where $\beta=(3c-1)/(2-2c)$ and $c_{1}^{\prime}$ and $c_{2}^{\prime}$ are coefficients fixed by the initial conditions. Using the asymptotic form of the Bessel functions for large real arguments this result can be simplified at either late times or large $k$ to
@@
The perturbations $\delta\phi$ can be treated as a separate radiation fluid distinct from the spatially-averaged scalar field $\bar{\phi}$. This is because at leading order, the only direct coupling between the two in the scalar field equation of motion occurs through the mass term $m^{2}(\bar{\phi})\delta\phi=\lambda^{2}M_{P}^{-2}V(\bar{\phi})\delta\phi$. As long as this term is negligible, as occurs for sub-horizon modes, then the two decouple. Importantly, this is the case irrespective of which component dominates the energy density of the scalar field. In particular, a consistent expansion of the exponential in Eq. (2.4) only requires that the inhomogeneities are small relative to the Planck scale $\lambda\delta\phi\ll M_{P}$, but they need not be small relative to the background scalar field or its time derivative.
@@
For sufficiently steep exponential potentials with $\lambda>2$, this cosmology has a well-known late-time solution given by the radiation tracker for which $c=1/2$. In this case, such a tracker solution can be made explicit by defining the kinetic and potential energy fraction of the full scalar field as Copeland:1997et
@@
$x=\langle\tilde{x}\rangle=\frac{\bar{\pi}}{\sqrt{6}HM_{P}},\ \ \ \ \ \ y=\sqrt{\frac{V(\bar{\phi})}{3}}\frac{1}{HM_{P}}=\langle\tilde{y}\rangle[1+\tfrac{\lambda^{2}}{8M_{P}^{2}}\langle\delta\phi^{2}\rangle+\mathcal{O}(\langle\delta\phi^{4}\rangle)].$ (2.21)
@@
$x^{\prime}(N)$ $=-3x+\lambda\sqrt{\frac{3}{2}}y^{2}+2x[1+\tfrac{1}{2}x^{2}-y^{2}],$ (2.22)
$y^{\prime}(N)$ $=-\lambda\sqrt{\frac{3}{2}}xy+2y[1+\tfrac{1}{2}x^{2}-y^{2}],$ (2.23)
@@
where primes represent derivatives with respect to $N$. For sufficiently steep exponential potentials with $\lambda>2$ this has the standard attractor radiation tracker solution
$x^{2}=\frac{8}{3\lambda^{2}},\ \ \ \ \ \ \ \ y^{2}=\frac{4}{3\lambda^{2}}.$ (2.26)
@@ 2.2 Non-linearities and Non-Exponential Potentials
It is useful to contextualize the case of exponential potentials by considering the case of more general potentials $V(\phi)$ (some of these results have recently been pointed out by Eroncel:2025qlk ). In that case the scalar field equation of motion can be written as
@@
$\delta\phi_{\boldsymbol{k}}(t)=\frac{c_{1}}{a^{3/2}\sqrt{2\omega}}\sin\bigg{(}\int^{t}dt^{\prime}\omega(t^{\prime})\bigg{)}+\frac{c_{2}}{a^{3/2}\sqrt{2\omega}}\cos\bigg{(}\int^{t}dt^{\prime}\omega(t^{\prime})\bigg{)},$ (2.28)
@@
In this work, we have primarily focused on the case of exponential potentials, where the potential factorizes into two exponentials $e^{-\lambda\bar{\phi}/M_{P}}e^{-\lambda\delta\phi/M_{P}}$. The equation for the perturbations can then be expanded out as long as $\lambda\delta\phi\ll M_{P}$, giving a linear equation, with higher order terms suppressed by factors of $\lambda\delta\phi/M_{P}$. Additionally, the approximation becomes more accurate with time as long as the amplitude of the perturbations decreases with time. This is true irrespective of whether $\delta\phi$ dominates the energy density of the universe or not. For example, the transition from a kinating epoch, where the energy density is dominated by the kinetic energy in the background field $\dot{\bar{\phi}}^{2}/2\gg V(\phi)$, to one where the universe is dominated by the perturbations, is perfectly linear in terms of the equation of motion for the perturbations $\delta\phi$.
In this work we have neglected metric backreaction from the scalar field, which has been shown to be negligible for adiabatic modes in linear perturbation theory during kination and radiation domination Eroncel:2025bcb . The same argument can be extended to the self-tracker to show that the backreaction is negligible in that case as well. [^2] The only difference from the radiation dominated case of Eroncel:2025bcb is the appearance of an additional $2\lambda\bar{V}\Phi$ term on the right hand side of the scalar field equation of motion (Eq. C.4 in Eroncel:2025bcb ), which for subhorizon modes scales like $t^{-3}$ and so becomes negligible with time. It would be interesting to consider non-adiabatic modes, but this is left for future work. If perturbations are large, the metric backreaction can become non-linear and have a significant impact on the evolution of the scalar field modes (see the recent paper KingsGroup ). This can occur when the energy density in some region is substantially larger than the background $\delta\rho/\bar{\rho}\gtrsim 1$ for long enough that the region decouples from the Hubble flow Aurrekoetxea:2024mdy . In that case a full numerical relativistic analysis is required, with these regions possibly giving rise to black holes. The case of a massless scalar field with an initial Gaussian profile has been studied in numerical relativity Yoo:2018pda . This system collapses to a black hole for sufficiently large initial amplitudes, but would otherwise transition into a radiation dominated universe if collapse is avoided, which corresponds in our case to perturbation domination with a negligible potential.
@@ 2.3 Perturbation-Dominated Eras
$\frac{1}{2}\bar{m}^{2}\langle\delta\phi^{2}\rangle=\bar{V}\frac{\lambda^{2}}{M_{P}^{2}}\int\frac{d^{3}\boldsymbol{k}}{(2\pi)^{3}}|\delta\phi_{\boldsymbol{k}}|^{2}\leq\bar{V}\bigg{(}\frac{\lambda}{M_{P}}\int\frac{d^{3}\boldsymbol{k}}{(2\pi)^{3}}|\delta\phi_{\boldsymbol{k}}|\bigg{)}^{2}\ll\bar{V},$ (2.29)
@@
$\frac{\lambda}{M_{P}}\int\frac{d^{3}\boldsymbol{k}}{(2\pi)^{3}}|\delta\phi_{\boldsymbol{k}}|\ll 1,$ (2.30)
which is a marginally stronger requirement than $\lambda\delta\phi/M_{P}\ll 1$, needed to be able to expand out the exponential to leading order. Since perturbation domination requires the energy density contribution from the perturbations to be much greater than that of the background potential $\bar{V}$, perturbation domination with small amplitudes cannot behave like a matter dominated era. [^4] For other potentials, this need not be the case. For example, for quadratic potentials where the background scalar field is at its minimum $\bar{V}=0$, one can have a matter dominated era driven by the energy density coming from the perturbations around the minimum.
@@
$3H|\delta\dot{\phi}_{\boldsymbol{k}}|\gg 3H\bar{m}|\delta\phi_{\boldsymbol{k}}|\gg\frac{1}{\lambda}\bar{m}^{2}|\delta\phi_{\boldsymbol{k}}|\sim\bar{m}^{2}|\delta\phi_{\boldsymbol{k}}|,$ (2.33)
where for the second inequality we have used that during perturbation domination $H^{2}\gg\bar{V}/M_{P}^{2}=\bar{m}^{2}/\lambda^{2}$. Therefore, for $\lambda\sim\mathcal{O}(1)$, the mass term is much smaller than the Hubble term. We have thus shown that when the mass contribution to the energy density is much smaller than the gradient/kinetic terms, then the gradient/Hubble terms dominate over the mass term in the equation of motion, respectively. The equation therefore corresponds to that of a massless scalar field, implying that during perturbation domination with subhorizon modes, the universe behaves like a radiation era.
@@
$\delta\ddot{\phi}_{\boldsymbol{k}}+3H\delta\dot{\phi}_{\boldsymbol{k}}+\bigg{(}\frac{k^{2}}{a^{2}}+\frac{\lambda^{2}\bar{V}}{M_{P}^{2}}\bigg{)}\delta\phi_{\boldsymbol{k}}=0,$ (2.34)
which for $k/a<\lambda\sqrt{\bar{V}}/M_{P}$ behaves like a massive scalar field. In a cosmology with $H=c/t$, this can be solved to give
@@
where $\beta=(3c-1)/2$, and in the second step we used the asymptotic form of the Bessel functions for large $mt$. This solution behaves like a matter field with mass $m=\lambda\sqrt{\bar{V}}/M_{P}$. However, so long as the background scalar field can start rolling down the exponential, then such matter behavior would be short-lived.